OpenStudy (anonymous):

Write the equation of the horizontal line passing through the point (-1, -8). i'm not sure if I put the right answer.

OpenStudy (anonymous):

is it x=-1?

OpenStudy (anonymous):

And also.... Write the equation of the vertical line passing through the point (3, 1). wud it be y=1? or x=3?

OpenStudy (anonymous):

plz helpppp

OpenStudy (anonymous):

To find the equation, I would figure out the slope, than plug everything into the point-slope form to get the original equation (standard form). Point-slope form: y – y1 = m(x – x1) Standard form: y = mx + b

OpenStudy (anonymous):

sorry I'm just confused with this slope thing i don't know if the equation shud have a slope or what?

OpenStudy (anonymous):

Yes, it should.

OpenStudy (anonymous):

but this lesson had nothing to do with slopes at least by what i know

ganeshie8 (ganeshie8):